OpenStudy (buttermequeasy):

Consider the line that passes through the points (0, 4) and (-1, 3). Below are two different equations. Identify the true statement concerning both equations. Equation #1 Equation #2 y - 4 = 1(x - 0) y - 3 = 1(x - 1) Only equation #1 represents the line that passes through the two given points. Only equation #2 represents the line that passes through the two given points. Both equation #1 and equation #2 represent the line that passes through the two given points. Neither equation #1 nor equation #2 represents the line that passes through the two given points.

OpenStudy (buttermequeasy):

help please! needed urgently!

OpenStudy (anonymous):

Only equation #1 represents the line that passes through the two given points. Just replace the points in the equations and look at the relationships that are true. For (0,4) y-4=x -> 4-4=0 (ok) (-1,3) 3-1=-1 For (0,4) y-3=x-1 --> 4-1is different of 0-1 (-1,3) 3-3 is different of -1-1